OpenStudy (anonymous):

evaluate each finite series for the specified number of terms. 80-40+20-.....;n=8

OpenStudy (anonymous):

@whpalmer4 @eliassaab

OpenStudy (anonymous):

can you help me??? @eliassaab

OpenStudy (anonymous):

\[ 80-40+20-10+5-\frac{5}{2}+\frac{5}{4}-\frac{5}{8}=\frac{425}{8} \]

OpenStudy (anonymous):

thank you but how did you get that

OpenStudy (anonymous):

each term is -the term before it divided by 2

OpenStudy (anonymous):

oh i see thats cool thank you

OpenStudy (anonymous):

\[ a_1=80\\ a_n =- \frac {a_{n-1}}2 \]

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

You list them and you do the arithmetic

OpenStudy (anonymous):

YW